Make New Entry,
Make Followup Entry
User name Meziani
Log entry time 08:03:45 on November 22, 2009
Entry number 301767
00:00 Shift start
Target in Empty position
00:20 ask for raster off
Start the checking/recover procedure
BPMA x = -0.665, y = 0.02
BPMB x = -0.650 y = 0.03
00:30 Harp scan completed waiting for MCC to log the results.
00:34 Harp scan results good see Elog: 1499492
00:38 Moving to Beo target
00:40 Asked for 2uA with raster off/ Someone please check posted
instructions current to see the beam on BeO greater than 2 uA.
could not see beam on the Beo although everyother indication on beam
position was good.
00:50 Asked MCC for 4x4 raster size to check spot++ results.
01:02 Results of Spot++ good
01:04 ask MCC to mask target, moving to LD2 position.
01:12 target is in position asked MCC for 2uA, raster 4x4 (MCC units)
01:16 Rates checked against witness rates: Ok
Taking L-HRS T1 data with LD2 and 3x3mm^2 spot.
Spot++ results look good not hot walls
01:20 asking for a 6x6(MCC units) raster size
01:22 Raster size 6x6 and checking with spot++
01:27 Spot++ results good 4x4 mm^2 roughly with no hot regions on the edge.
01:28 turn the VDC HV to off (0.00V).
01:29 asking MCC to raise the current to 60uA (should be ok according to
01:31 asking MCC for 80 uA.
01:32 asking MCC for 100 uA
01:36 2 trips at 100 uA.
01:43 4 trips at 100 uA.
MCC called mentioning the lack of orbit lock for Compton
02:26 after seeing too many trips due to Compton ion chamber
I requested to lower the current to 95uA to have less trips.
Things seems to be runnig better at below 100 uA.
02:33 Megan called MCC and mentioned that the orbit lock was on. MCC
Time run#PVDIS run#HRS-L run#HRS-R # evts comment
00:00-00:59 14130 Pedestal
01:14 01:27 14131 5223 25672 29.8K/0.22M/0.52 Calib.
01:36-02:34 14132 5224 25673 100K/3.44M/3.21M Prod.
02:38-03:34 14133 5225 25674 100K/3.48M/3.25M Prod.
03:35-04:34 14134 5226 25675 100K/3.76M/3.54M Prod.
04:37-05;32 14135 5227 25676 100K/3.42M/3.2M Prod.
05:35-06:34 14136 5228 25677 100K/3.70M/3.2M Prod.
06:35-07:35 14137 5229 25678 100K/3.42M/3.20M Prod.
Call MCC to change the IHWP from the IN to the OUT position
07:36 IHWP is OUT
Ignore 1438, 5230 and 25679 (junk)
07:40 14139 5231 25680
8:00 am Shift Ended in middle of above runs